OpenStudy (anonymous):

Prove that for any constant a, |r|<1, then sum ar^k=a/(1-r) for k from 0 to infinity

OpenStudy (anonymous):

the sum of a geometric progression is a(1-r^n)/(1-r)

OpenStudy (anonymous):

Typing these proofs out in latex takes so long... X_X First we'll start with: \[a+ar+ar^2+ar^3+...\] \[= \lim_{n \rightarrow \infty}(a+ar+ar^2+ar^3+ ... +ar^n)\] Now we can use the formula for geometric progression: \[a+ar+ar^2+ar^3+...+ar^n = a(1 - r^{n+1})/(1 - r)\] and then we get: \[= \lim_{n \rightarrow \infty}(a(1 - r^{n+1})/(1 - r))\] \[r^{n+1} \rightarrow 0\space for\space|r| < 1\] \[\therefore\space=a/(1-r)\]

OpenStudy (anonymous):

as r < 1 r^inf -> 0 therefore you get a*1/(1-r)